OpenStudy (anonymous):

Calculate the volume of a cylinder that has a diameter of 8 inches and a height of 3 inches.

OpenStudy (diyadiya):

\[\text {Volume of a cylinder}=\pi \times radius^2 \times height\]

OpenStudy (anonymous):

48pi inch^2